### v1

# Counterfactual Visual Explanation via Causally-Guided Adversarial Steering

Abstract
Recent work on counterfactual visual explanations has contributed to making artificial intelligence models more explainable by providing visual perturbation to flip the prediction. However, these approaches neglect the causal relationships and the spurious correlations behind the image generation process, which often leads to unintended alterations in the counterfactual images and renders the explanations with limited quality. To address this challenge, we introduce a novel framework CECAS, which first leverages a causally-guided adversarial method to generate counterfactual explanations. It innovatively integrates a causal perspective to avoid unwanted perturbations on spurious factors in the counterfactuals. Extensive experiments demonstrate that our method outperforms existing state-of-the-art approaches across multiple benchmark datasets and ultimately achieves a balanced trade-off among various aspects of validity, sparsity, proximity, and realism.
1 Introduction
Despite the impressive performance of deep neural networks (DNNs) across domains such as language, vision, and control [1, 2, 3], explaining their predictions remains challenging due to their non-linear, overparameterized nature [4]. This challenge is particularly critical in safety-sensitive applications like healthcare, finance, and autonomous driving [5]. As a result, explainable AI (XAI) has gained momentum [6], with counterfactual explanations (CEs) [7] emerging as a powerful approach. CEs aim to answer: What minimal change to an input would alter the model’s prediction? They are valued for producing intuitive and preferably actionable insights that can support decision auditing and planning. As a powerful XAI tool [7], CE has been particularly flourishing in computer vision [8, 9, 10, 11]. Unlike attribution-based methods that highlight pixels associated with the output, counterfactuals in images simulate realistic alternative scenarios, allowing us to ask: What would the image need to look like for the model to decide differently? For example, modifying a few facial features to change a classification from “neutral” to “happy” provides insight into the model’s decision boundary. Successful visual counterfactuals not only satisfy conventional CE criteria such as being valid (flip the prediction), proximal (close to the original image), and sparse (modify as few features as possible) [12, 13], but also are semantically meaningful and perceptually natural, avoiding unrealistic artifacts. These methods offer intuitive and human-understandable explanations, making them especially valuable for analyzing high-dimensional visual models.
While counterfactual visual explanation has seen rapid progress in recent years [8, 9, 10, 11], several critical aspects of the problem remain overlooked. A key limitation lies in the lack of a principled understanding of the underlying causal mechanisms in the data, as well as the true target of explanation. Most existing methods [14, 15, 16] optimize directly over the entire input feature space in search of perturbations, which often leads to changes in spurious factors that are correlated with the label but not causally related. While such perturbations may achieve a prediction flip, they rarely reflect meaningful or actionable changes in the real world and can even produce misleading or harmful outcomes. For example, in autonomous driving scenarios, changing the shape of a traffic sign (spurious factor) to alter a “stop” vs. “go” prediction is far less informative than changing the content of the sign, which is causally relevant to the model’s decision. More importantly, conventional CE strategies often result in low-quality and unrealistic images. The pixel space of images is vast, and unconstrained perturbations can introduce noise or artifacts that are visually implausible or semantically incoherent. Recent advances [9, 11] have begun to address this by leveraging generative models, particularly diffusion-based methods, which operate in semantically meaningful latent spaces and are better equipped to produce realistic counterfactuals. These approaches benefit from the ability of diffusion models to suppress high-frequency noise and structure perturbations more coherently. However, most current methods still rely heavily on predefined loss functions, such as $l_{1}$ loss constraining the difference between original and counterfactual images, to enforce sparsity and proximity. While effective at a global level, these constraints often lead to unwanted changes in local artifacts such as color shifts, shape distortions, or the emergence of spurious attributes [17]. This results in a persistent trade-off between validity, proximity, sparsity, and realism, making it difficult to generate counterfactuals that are both high-quality and causally meaningful. To the best of our knowledge, no existing method well addresses these challenges, underscoring the need for a new approach that incorporates causal reasoning into the generation of counterfactual visual explanations.
To address the aforementioned challenges, we propose a novel framework, Counterfactual Explanation via Causally Adversarial Steering (CECAS), which leverages a causality-guided adversarial method to generate counterfactual explanations that are more semantically faithful and causally grounded. Adversarial attacks have been an effective tool to generate CEs due to their common aim to minimally alter an input to change a model’s prediction [18]. Uniquely, our approach integrates causality into the adversarial objective: While the adversarial component ensures prediction flips (validity), the causal guidance directs perturbations toward features that are causally relevant to the label. This promotes sparsity, proximity, and realism, by avoiding changes to spurious or irrelevant features. As a result, our method produces high-quality, semantically meaningful counterfactual images that align with real-world interventions. We conclude our contributions as follows:
  • We provide a new perspective on the problem of counterfactual visual explanation by highlighting the critical role of causality. We show how causality can improve the quality of counterfactuals by avoiding spurious perturbations.
  • We propose a novel framework with a causally guided adversarial approach to generate counterfactual visual explanations. Our method steers perturbations toward causally relevant features, enabling the generation of valid, sparse, proximal, and realistic counterfactual images.
  • We conduct extensive experiments to evaluate our framework on multiple benchmark datasets. Our approach outperforms current state-of-the-art methods across various metrics both quantitatively and qualitatively.
2 Related Work
Explainable AI (XAI) XAI refers to a set of techniques aimed at enhancing the transparency and explainability of AI models—making their decisions and internal mechanisms more understandable to humans [5]. The field is commonly divided into two major categories: ad-hoc and post-hoc methods. The key difference between them lies in the stage at which they are implemented [19]. The former is often deployed during the training phase and advocates the design of inherently interpretable architectures [20, 21, 22, 23, 24]. In contrast, the latter focuses on understanding and explaining the trained models without modifying their internal structure [25, 26, 27]. The post-hoc category includes global and local explanations, distinguished by the scope of the explanations they provide [5]. Global explanation provides a comprehensive description of the model [28], while the latter focuses on explaining the model’s prediction for a specific instance [29]. The studied problem of our paper belongs to local explanations. This line of work mainly focuses on the following three directions: saliency map [30, 31, 32], concept attribution [33, 34, 35], and model distillation [36, 37].
Counterfactual Explanations As a branch of post-hoc explanations, CE has emerged as a highly influential approach in recent years [18], especially in computer vision [11]. Some methods generate counterfactual samples by locating key regions that distinguish two contrasting images and substituting the relevant region from one image into the other [38, 39]. Other methods utilize the gradient of the input image with respect to the target label while also leveraging robust models to generate human-recognizable modifications [40, 41]. Additional approaches, including prototype-based algorithms [42] and invertible convolutional networks [43], have also shown promising abilities in generating counterfactual examples. Generative techniques are inherently well-suited for CE. With their rapid development, an increasing number of related works have emerged, which can be broadly categorized into conditional [44, 45] and unconditional approaches [46, 10]. Diffusion models (DDPM) are cutting-edge generative models. Jeanneret et al. [9] and [11] both utilize DDPM to generate counterfactual images. The difference lies in that the former modifies the sampling process of the DDPM, while the latter employs DDPM as a regularizer.
Adversarial Attacks Adversarial attacks and CEs share a similar objective: flipping the model’s prediction, which has led some studies to focus on exploring the relationship between them [47, 48, 18]. However, the key difference lies in the fact that adversarial attacks on undefended models often result in imperceptible noise rather than semantically meaningful changes. Adversarial attacks include white-box and black-box settings. In white-box attacks, the attacker has access to the model’s architecture and data [49]. It typically generates adversarial examples by utilizing the gradient of the loss function with respect to the input [50, 51, 52]. In contrast, black-box attacks [53, 54] restrict the attacker to accessing only samples from an oracle or pairs of inputs and corresponding outputs.
Adversarial Defense The development of adversarial attacks has, in turn, stimulated significant progress in adversarial defense. Recent works can be primarily divided into two categories: certified defense and empirical defense. Certified defense [55, 56, 57] offer provable robustness by ensuring that the model’s predictions remain stable under perturbations constrained within a specified norm ball. However, empirical defense [50, 51, 58], particularly adversarial training [51], is still highly effective. It typically generates adversarial samples by using an adversarial attack first and then incorporates such samples into the training process. It is worth noting that recent work has begun to integrate causality into adversarial defense strategies [59, 60, 61, 62], sharing a common objective with our work.
3 Preliminaries
Problem Definition We first consider a classification task with $x$ representing an input image and $y$ standing for the corresponding label, where $x\in\mathcal{X},y\in\mathcal{Y}:=\{1,...,K\}$, $\mathcal{X}$ and $\mathcal{Y}$ represent the image space and the label space, respectively. A DNN classifier trained for this task can be denoted by $f:\mathcal{X}\rightarrow\mathcal{Y}$. Then the counterfactual visual explanation can be defined as the task of designing a semantically meaningful, perceptible yet minimal perturbation $\delta$ that, when added to an input image $x$, generates a counterfactual image $x^{\prime}:=x+\delta$ and flips the classifier’s prediction $\hat{Y}$ from the original prediction $y$ to a target label $y^{\prime}$.
Figure 1: Causal graph of the natural and counterfactual images generation process. Each node means a variable, where blue nodes are involved in the generation of natural images, green nodes contribute to the generation of counterfactual images, and two-colored nodes participate in both.
Causal View on Image Generation Figure 1 is a well-recognized causal graph $\mathcal{G}$ [62] that can represent the generation processes of both the natural and counterfactual images. As shown, $C$ denotes the causal factors that directly determine the label $Y$ and predicted label $\hat{Y}$. $S$ denotes the spurious factors, which are independent of $C$, as reflected in the $v$-structure $C\rightarrow{X}\leftarrow{S}$. $X$ is the natural image, and $\Delta$ represents the semantically meaningful perturbation; together, they form the counterfactual image $X^{\prime}$. $\theta$ is the set of parameters of the DNN. It is evident from the graph that the perturbation $\Delta$ is determined by $\hat{Y}$, $X$ and $\theta$. The inclusion of $\theta$ is due to the use of white-box attacks, such as projected gradient descent (PGD). Note that $S$ can also have an influence on $Y$ because of the spurious paths in $\mathcal{G}$. When $X$ is given, there is a spurious path ${Y}\leftarrow{C}\rightarrow{X}\leftarrow{S}$ associate $Y$ and $S$.
Denoising Diffusion Probabilistic Models DDPM [63] can filter out the noise introduced by traditional adversarial attacks without changing the classifier’s weights. It relies on two Markov chains, corresponding to the forward and reverse processes, respectively. The forward process gradually adds Gaussian noise to the data at the state of $t$ to generate data at the next state $t+1$. In practice, the data at time step $t$, denoted as $x_{t}$ can be directly obtained from the clean data $x_{0}$ through
$x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\epsilon,\quad\epsilon\sim\mathcal{N}(0,I),$ (1)
where $\bar{\alpha}_{t}:=\prod_{s=1}^{t}\alpha_{s}$, ${\alpha}_{t}:=1-{\beta}_{t}$ and $\{\beta_{t}\}_{t=1}^{T}$ is the variance schedule. The reverse process removes the noise from $x_{t}$ to get a less noisy $x_{t-1}$ through the following equation:
${x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left({x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\,\bm{\epsilon}_{\rho}({x}_{t},t)\right)+\sigma_{t}\epsilon,$ (2)
where ${\sigma_{t}}^{2}={\beta}_{t}$, $\bm{\epsilon}_{\rho}$ is a function approximator intended to predict $\epsilon$ from $x_{t}$.
Projected Gradient Descent Attack PGD is a white-box attack that leverages the gradient of the loss function $L$ with respect to the input image $x$ to iteratively modify the image in a direction that maximally confuses the classifier. PGD operates through multiple iterations, and after each iteration $\tau$, the perturbed image $x_{\tau}$ is projected back onto an $l_{\infty}$-norm ball centered at the original input, with a predefined radius. This process can be summarized as follows:
$x_{\tau+1}=\Pi_{x+\mathcal{S}}\left(x_{\tau}+\alpha\,\text{sgn}\left(\nabla_{x}L(\theta,f(x_{\tau}),y)\right)\right),$ (3)
where $\Pi$ denotes the projection operator, $\mathcal{S}$ defines the radius of the $l_{\infty}$-norm ball, $\alpha$ is the step size, sgn is the sign function, and $\theta$ represents the parameters of the target model.
4 Methodology
Figure 2: CECAS Framework: The upper part of the figure illustrates the first stage of the framework, which involves generating the initial CE images. Here, $x$ denotes the original image, $x^{\prime}$ represents the causally-guided CE. The lower part of the figure depicts the second stage of the framework, namely post-refinement. Here, we refine $x^{\prime}$ with a precise pixel-level mask $m$ (computed based on the difference between $x$ and a $l_{1}$-based CE $x^{\prime\prime}$), subsequently used in conjunction with image inpainting techniques to produce the high-quality final CE image.
We propose the CECAS framework, which consists of two key components, as shown in Figure 2. The first component is the initial generation of CE images, achieved through a novel causally guided adversarial method. The second component is the causal refinement of the generated counterfactuals, where we adopt an image inpainting approach to better preserve the details of the unmodified regions of the original image. We detail them in the following sections.
4.1 Initial Causally-Guided Counterfactual Visual Explanation Generation
The first and central component of the proposed framework is a causally guided adversarial method to generate counterfactual visual explanations. As previously mentioned, adversarial attacks and CE share a common goal: altering a model’s prediction through minimal input perturbations. Inspired by prior work bridging the two domain [18, 9, 64], we adopt adversarial strategies—such as PGD attacks—to generate image-based counterfactuals. The procedure is outlined as follows:
$x_{\tau+1}=\Pi_{x+\mathcal{S}}\left(x_{\tau}-\alpha\,\text{sgn}\left(\nabla_{x}L(f_{\theta}(x_{\tau}),y^{\prime})\right)\right),$ (4)
where the subscript $\tau$ denotes the iteration in PGD, $L$ is the classification loss, typically the cross-entropy loss, and $y^{\prime}$ is the target label. Upon completing $T$ iterations, the resulting $x_{T}$ is taken as the counterfactual image we aim to generate. It is worth noting that, in form, Equation (4) differs from traditional PGD attacks. Here, the gradient is subtracted from the perturbed image rather than added. This is because the objective is not to mislead the classifier into making an arbitrary incorrect prediction, but rather to guide it toward a specific target label $y^{\prime}$. This allows CEs to be generated with the gradient-based perturbation $\delta$ added on the original image.
However, relying solely on the above adversarial attack-based method is often overly idealized, as these methods typically impose no constraints on spurious factors. As discussed in the Introduction, although the explained model may rely on such factors during prediction, perturbing them often does not yield meaningful or actionable explanations. Furthermore, altering these spurious factors can introduce artifacts or noise, which severely compromises the core goals of CE of providing realistic and constructive explanations.
To address this issue, it is essential to constrain changes to spurious factors. Given the central role of gradients in the above CE generation process, we aim to identify a new gradient direction—one that enables effective modification of the image while preserving spurious factors unchanged. Since our goal is to explain the classifier $f$, we are not allowed to make any modifications to it. Considering the architecture of the explained model $f$, a DNN commonly viewed—following [65]— as a combination of a feature extraction module and a linear classifier module, we learn to capture the spurious factors out of the representations $R$ learned by the feature extraction module (details of how to capture $S$ from $R$ are given in the final part of this subsection). With the captured $S$, we use the original data to train an auxiliary classifier $g$ which predicts the label $Y$ based on the spurious factor. More specifically, after training, $g$ can make predictions as follows: $\hat{z}=g(s),\hat{z}^{\prime}=g(s^{\prime})$, where $\hat{z}$ denotes the logit predicted by $g$ from the spurious factor $s$ in the original image, while $\hat{z}^{\prime}$ is the logit predicted by $g$ from the spurious factor $s^{\prime}$ in the modified image. We can consider $z$ to capture the spurious factors that are most relevant to the label $Y$—that is, the spurious features most likely to be unintentionally altered during counterfactual generation. On top of this, we introduce a penalty loss $L_{S}$ that encourages these spurious factors to remain invariant during the counterfactual generation process, and thereby we compute the gradient of the penalty loss with respect to the image and incorporate its direction into the guidance for image modification:
$x_{\tau+1}=\Pi_{x+\mathcal{S}}\left(x_{\tau}-\alpha\,\text{sgn}\left(\nabla_{x}L(f_{\theta}(x_{\tau}),y^{\prime})+\nabla_{x}L_{S}(z_{\tau},z_{\tau}^{\prime})\right)\right),$ (5)
where $L_{S}$ can typically be implemented as the soft cross-entropy loss, which takes two logits as input. The primary purpose of this loss is to align the spurious distributions between the natural and counterfactual representations by penalizing the deviation of spurious factors in the counterfactuals from those in the original inputs. In summary, the generation of counterfactual images involves modifying the images via adversarial attack, while preserving the spurious factors through a causal perspective.
Furthermore, inspired by the effectiveness of diffusion models in enhancing CE image quality [9, 64], we integrate diffusion models to filter out high-frequency, semantically meaningless noise introduced by adversarial attacks. Specifically, we apply the forward process of DDPM to add Gaussian noise to the image first, and then use the reverse process to denoise it before feeding it into the classifier for loss computation. In this work, we define the joint classifier $F$ as the combination of DDPM and the base classifier $f$. The generation of counterfactual images can be simplified and reformulated in an optimization form:
$\arg\min_{x^{\prime}}L(F(x^{\prime});y^{\prime})+\alpha L_{S}(z,z^{\prime}),$ (6)
where $\alpha$ is the hyperparameter that controls the strength of $L_{S}$.
The only remaining problem is how to obtain the spurious factor $S$. Fortunately, numerous recent works have explored disentangling causal factors from spurious ones [66, 62, 67]. Building on the causal model illustrated in Figure 1, we adopt the widely used assumption that the causal factor $C$ is independent of the spurious factor $S$. Leveraging this independence, we employ a subspace projection technique to project $C$ and $S$ to orthogonal subspaces and thereby extract the spurious factor $S$, as proposed in Zhang et al. [62], this technique is supported by solid theoretical foundations. Notably, our framework is not limited to this specific implementation and can be extended or replaced by other effective strategies for extracting $S$.
4.2 Causally-Guided Post Refinement
After generating the initial counterfactual images, a key challenge lies in ensuring sparsity and proximity. This is particularly difficult because the $L_{S}$ constraint operates at the concept level and often becomes “soft” when translated to pixel space, making it insufficient to enforce strict pixel-level consistency. While we have discussed the limitations of the $\ell_{1}$ loss, its advantages should not be overlooked—particularly its ability to identify precise pixel-level changes. Therefore, we leverage this property of $\ell_{1}$ loss to roughly localize pixels in the image that require modification, and then combine it with our causal constraint $L_{S}$ to avoid unnecessary changes to spurious factors. This supports our goal of causally guided post-refinement, where we preserve the counterfactual changes in $x^{\prime}$ introduced by causal adversarial steering, while restoring the unmodified regions with the corresponding pixels from the original image $x$. Specifically, we first generate a perturbed image using an adversarial attack combined with $l_{1}$ loss as shown in the following equation:
$\arg\min_{x^{\prime\prime}}L(F(x^{\prime\prime});y^{\prime})+\lambda||x^{\prime\prime}-x||_{1},$ (7)
where $x^{\prime\prime}$ is the perturbed image, $\lambda$ is a hyperparameter. Then a binary mask is constructed as:
$m_{i,j}=\mathds{1}\left[\max_{(u,v)\in\mathcal{N}_{i,j}}\left(\frac{1}{M}\sum_{ch=1}^{CH}\left|x^{\prime\prime}(ch,u,v)-x(ch,u,v)\right|\right)>\gamma\right],$ (8)
where $m_{i,j}$ is the binary mask value at position $(i,j)$, $x$ is the original image, $ch$ is the channel index, $(u,v)$ is the position index in an image, $CH$ is the number of channels, $M$ is the maximum value after the summation, $\mathcal{N}_{i,j}$ is a $d\times{d}$ neighborhood window centered at position $(i,j)$, $\mathds{1}$ is the indicator function and $\gamma$ is a predefined threshold. We apply the dilation operation to adjust the size of the modification region to a reasonable extent, allowing for as many plausible changes as possible.
The mask indicates which pixels need to be modified through $l_{1}$ loss and which do not. Intuitively, we need to combine $x^{\prime}$ generated in the last subsection and the original image $x$ with the mask to achieve causal perturbation in CEs with well-preserved pixels in the rest parts. However, directly combining $x^{\prime}$ and $x$ using the mask is not feasible, as it would result in clearly visible region boundaries in the CE. In our work, we leverage the image inpainting technique [68] to tackle this problem. The inpainting process is shown as follows:
${x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left([(1-m){x}_{t}+mx^{0}_{t}]-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\,\bm{\epsilon}_{\rho}([(1-m){x}_{t}+mx^{0}_{t}],t)\right)+\sigma_{t}\epsilon,$ (9)
where $m$ is the binary mask, $x_{t}$ denotes the image being refined at the noise level $t$, and $x^{0}_{t}$ is the original image at the noise level $t$. After a pre-defined $T$ denoising steps, we can obtain the final counterfactual visual explanations $x^{\prime}$ from the previous initial $x^{\prime}$ we gained from Section 4.1.
5 Experiments
In this section, we present extensive experiments to evaluate our proposed framework. We first compare our method against state-of-the-art baselines through both qualitative and quantitative analyses, including widely used metrics and case studies. Beyond traditional metrics, we further assess the quality of the generated counterfactuals using vision-language models (VLMs) as human-like evaluators. Additionally, we perform parameter studies to better understand the impact of different hyperparameter values in our method.
5.1 Experimental Setup
Datasets We conducted extensive experiments on 8 datasets in total. We first evaluate our framework on the CelebA dataset [69], using images of size 128$\times$128 and focusing on two attributes which constitute two sub-datasets: “Smile” and “Age”. The CelebA-HQ [70] dataset is used with the same attributes and is also constructed into two sub-datasets, but with images of size 256$\times$256. The other datasets include the traffic scene dataset BDD-OIA [71], with images of size 512$\times$256, and three sub-datasets extracted from ImageNet [72], each containing two classes. These three sub-datasets are: Cougar vs. Cheetah, Sorrel vs. Zebra, and Egyptian Cat vs. Persian Cat. All experiments were conducted on NVIDIA RTX 4090 and RTX A6000 GPUs.
Baselines We selected four recent and state-of-the-art methods as our baselines: STEEX [8], DiVE [10], DiME [9], and ACE [11]. Among them, DiVE includes an extended variant called DiVE100, in which the number of optimization steps is reduced to 100.
Models Weights and Architectures To ensure a fair comparison, we use the same pre-trained DDPM weights as those used in the baselines for each dataset. For ImageNet, we use the ResNet-50 model pre-trained by PyTorch [73] as the classifier. For the other datasets, we use DenseNet-121 as the classifier and adopt the pre-trained weights provided by the baselines. Additional training details, including hyperparameter settings, are provided in Appendix A.
Evaluation Metrics We evaluate our method and the baselines across multiple dimensions, each assessed using several metrics. The first evaluation dimension is validity, which we measure using Flip Rate (FR) and Counterfactual Transition (COUT). FR denotes the proportion of samples for which the counterfactual successfully causes the classifier to change its prediction. COUT measures the transition scores between the original and counterfactual images, with values ranging from $-1$ to $1$. The second evaluation dimension focuses on sparsity and proximity, which we assess using Learned Perceptual Image Patch Similarity (LPIPS) [74], Correlation Difference (CD) [9], and SimSiam Similarity (S3) [75]. LPIPS measures the perceptual similarity between two images based on human visual perception. CD computes the average number of non-target attributes that are unintentionally altered across all samples. S3 evaluates the cosine similarity between the original and counterfactual images using SimSiam as the encoding network. The third evaluation dimension is realism, which we measure using Fréchet Inception Distance (FID) [76] and split FID (sFID) [11]. Since counterfactual images typically share most pixels with the original images, FID tends to be underestimated and may fail to reflect their true differences. To address this issue, sFID is proposed. It splits both the original and counterfactual image sets into two subsets, computes FID in a cross-wise manner, and repeats this process ten times to obtain an average score. All results reported in this paper are averaged over multiple experimental runs. We omit standard deviations as they are negligible compared to the corresponding mean values.
Table 1: Evaluation Results on CelebA: Smile.
Methods FR $\uparrow$ COUT $\uparrow$ LIPIPS $\downarrow$ S3 $\uparrow$ CD $\downarrow$ FID $\downarrow$ sFID $\downarrow$
DiVE - - - - - 29.4 -
DiVE100 - - - - 2.34 36.8 -
STEEX - - - - - 10.2 -
DiME 97.60% -0.0451 0.0896 0.9158 1.99 12.3 25.48
ACE 99.05% 0.6547 0.0273 0.9822 2.05 4.2 26.38
Ours 99.40% 0.7325 0.0138 0.9834 2.09 3.8 25.70
Table 2: Evaluation Results on CelebA: Age
Methods FR $\uparrow$ COUT $\uparrow$ LIPIPS $\downarrow$ S3 $\uparrow$ CD $\downarrow$ FID $\downarrow$ sFID $\downarrow$
DiVE - - - - - 33.8 -
DiVE100 - - - - - 39.9 -
STEEX - - - - - 11.8 -
DiME 73.10% 0.0381 0.0938 0.9158 3.81 20.2 34.37
ACE 99.00% 0.5441 0.0326 0.9843 3.37 4.3 26.47
Ours 99.95% 0.6908 0.0167 0.9830 2.98 4.6 26.26
Table 3: Evaluation results on BDD-OIA.
Methods FR $\uparrow$ COUT $\uparrow$ LIPIPS $\downarrow$ S3 $\uparrow$ FID $\downarrow$ sFID $\downarrow$
DiME 60.50% -0.4865 0.1119 0.9638 62.30 225.17
ACE 99.90% 0.7278 0.0318 0.9956 9.50 91.70
Ours 99.90% 0.8240 0.0150 0.9952 13.85 92.48
Table 4: Evaluation results on Imagenet.
Datasets Methods FR $\uparrow$ COUT $\uparrow$ LIPIPS $\downarrow$ S3 $\uparrow$ FID $\downarrow$ sFID $\downarrow$
Sorrel- Zebra ACE 76.67% 0.0967 0.1670 0.9285 77.76 124.66
Ours 90.00% 0.5442 0.1714 0.9367 74.58 114.66
Cougar- Cheetah ACE 76.67% 0.2059 0.1809 0.9625 108.67 147.79
Ours 93.33% 0.3390 0.1855 0.9693 105.49 146.04
Persian Cat- Egyptian Cat ACE 80.00% 0.2326 0.1531 0.8939 139.65 227.07
Ours 83.33% 0.3424 0.1543 0.8759 134.57 213.21
5.2 Quantitative Analysis
In this section, we conduct a quantitative analysis of all datasets with results summarized in Table 1 through Table 6. We highlight the best performance in bold and the second-best in italics. The results for the baselines DiVE, DiVE100, and STEEX, we extract the results from the original papers. For the data not reported in the original papers, we leave the corresponding entries blank in the tables.
Tables 1 and  2 show the results on the CelebA dataset. As observed, our method achieves the best performance on 5 out of the 7 metrics and ranks second on the remaining two. More specifically, our method achieves the highest FR and COUT scores, while also attaining the lowest LPIPS and S3 values, along with relatively low FID and sFID scores. This indicates that our method can achieve the highest flip rate with minimal modifications. In other words, it attains the highest validity while preserving both sparsity and proximity. Due to the similarity between the CelebA-HQ and CelebA datasets, we include the results for CelebA-HQ in Appendix B. Table 3 presents the results on the BDD-OIA dataset, which is the most complex among all, containing rich traffic-related elements. Due to space constraints, we omit the results of DiVE and STEEX on this dataset with a similar observation in other benchmarks. Our method did not improve the FR on this dataset because the FR is already near its upper bound. However, it still achieves the highest COUT score without incurring significant degradation in FID and sFID. This suggests that CECAS is capable of identifying more effective counterfactual transition directions, leading to increased classifier confidence in the target label prediction. Table 6 reports the results on the three ImageNet subsets. Due to the page limit, we focus on the comparison between our framework and the most competitive baseline ACE. Apart from marginally lower performance in LPIPS and S3 on the final subset, our method significantly outperforms ACE on all other metrics.
Figure 3: Qualitative results on CelebA dataset. Red circles highlight unintended changes produced by ACE, while green circles indicate the corresponding improvements made by our method.
Figure 4: Qualitative results on BDD-OIA dataset. Green circles highlight the counterfactual changes. The right column for each transition direction shows a zoomed-in version of the left column.
Figure 5: Qualitative results on ImageNet dataset. Green circles highlight the counterfactual changes.
5.3 Qualitative Analysis
In this section, we perform case studies to straightforwardly evaluate the quality of the generated counterfactual images from a human perceptual perspective. Due to the complexity of the BDD-OIA and ImageNet datasets, it is difficult to discern the differences between counterfactual images generated by our method and those produced by other methods through visual inspection. Therefore, for these two datasets, we only present the results of our method. For the other datasets, we compare the counterfactual images generated by our approach with those produced by the best-performing method, ACE, to demonstrate the visual superiority of our method. As shown in Figure 3, ACE introduces inexplicable color and shape distortions, as well as unintended attributes in the generated images. For example, in the transformation from “young” to “old,” some female samples are incorrectly rendered with dark patches on the chin—an artifact caused by the spurious correlation between beard and aging in male samples. This results in unmeaningful CEs. In contrast, our method effectively mitigates such artifacts by avoiding reliance on spurious cues. Figures 4 and 5 present the results on complicated datasets BDD-OIA and ImageNet, respectively. The green circles in the images highlight the necessary and human-perceptible changes made to achieve counterfactual visual explanations. We include the results for CelebA-HQ in Appendix B for the same reason. More detailed results for the other datasets are also provided in Appendix B.
5.4 VLM Evaluations
Table 5: Performance comparison across datasets.
Datasets DiME ACE Ours
CelebA Smile 7.49 8.14 8.24
CelebA Age 7.06 7.13 7.57
CelebA-HQ Smile 6.18 8.19 8.55
CelebA-HQ Age 3.95 5.76 6.19
BDD-OIA 4.07 7.86 7.77
Sorrel-Zebra - 5.73 6.37
Cougar-Cheetah - 7.14 7.61
Persian Cat-
Egyptian Cat
- 6.93 7.63
Vision-Language Models (VLMs), through large-scale pretraining on image–text pairs, demonstrate strong reasoning capabilities in tasks such as visual question answering and image captioning [77]. Overcoming the limitation of traditional CE metrics that only provide shallow numerical evaluations and fail to capture the semantic meaning of images, VLMs provide more human-like evaluations by leveraging their rich prior knowledge in pertaining [78]. To better suit the specific nature of counterfactual image generation, we carefully design unique prompts with evaluation criteria covering validity, sparsity, proximity, and realism. The prompt design for each dataset can be found in Appendix C. Notably, sparsity in our prompt is partially assessed from a causal perspective, examining whether the modified regions in the image are necessary or causally relevant, while ensuring that unrelated parts remain unchanged. In the context of our task, we utilize GPT-4o to assign scores ranging from 1 to 10 for each of the four evaluation dimensions. Given that these dimensions hold equal importance in counterfactual generation, the final score is computed as the average across all four. As shown in Table 6, our method outperforms the baselines in 7 out of the 8 datasets. Combined with the powerful visual and textual reasoning capabilities of VLMs, we have strong reason to believe that the counterfactual images generated by our method offer the highest level of interpretability.
5.5 Parameters Study
Figure 6: Line plots of FR and FID with respect to varying values of $\alpha$.
In this section, we analyze the impact of the hyperparameter $\alpha$ on counterfactual image generation. We present results on the CelebA: Smile dataset, while results on the other datasets are provided in Appendix D. As shown in Figure 6, FID initially decreases and then increases as $\alpha$ increases. This is because, in the early stages of increasing $\alpha$, the constraint on changes to spurious factors becomes stronger, reducing unintended modifications and thereby making the generated counterfactual images closer to real ones. However, as $\alpha$ continues to increase, the counterfactual generation process begins to overemphasize the alignment of spurious distributions. As a result, regions that originally required slight modifications become overly constrained, making the generated images more unreal. Consequently, the FID score gradually increases again. Meanwhile, the increase in $\alpha$ leads to a decline in FR, this is reasonable since when $\alpha=0$, the flip rate is highest as the optimization focuses solely on flipping the prediction. However, this comes at the cost of poor counterfactual quality, often producing unrealistic images. Therefore, it is crucial to strike a balance between flip effectiveness and visual realism. Fortunately, with a properly chosen $\alpha$, our method effectively achieves this balance, producing counterfactuals with both high flip rates and high image quality.
6 Conclusion
In this paper, we propose a novel framework, CECAS, which innovatively designs a causally-guided adversarial method to generate counterfactual visual explanations that constrain less informative changes on spurious factors. The framework can be applied to any explained DNN models. Experimental results demonstrate that our method consistently outperforms existing approaches in quantitative evaluation, qualitative analysis, and VLM-based assessment. CECAS achieves a better balance among the four key properties of counterfactual explanations: validity, sparsity, proximity, and realism. Therefore, we argue that the counterfactual explanations generated by our method offer a more causally grounded understanding of the classifier’s decision-making and prediction mechanisms.
Limitations While our method focuses on factors that are causally related to the label, it does not explicitly model more fine-grained causal relationships between more variables, which could further enhance consistency in counterfactual generation. Exploring such nuanced causal structures remains a valuable direction for future work. Additionally, developing techniques to automatically balance competing objectives would be important toward more adaptive and generalizable CE methods.
Appendix
A. Training Details
Data Preparation The eight datasets discussed in the main text are categorized into two types: binary datasets and multi-label datasets. CelebA Smile, CelebA Age, CelebA-HQ Smile, CelebA-HQ Age, and BDD-OIA belong to the first category, while Cougar vs. Cheetah, Sorrel vs. Zebra, and Egyptian Cat vs. Persian Cat fall into the latter. For binary-label datasets, if the original predicted label is $y$ ($y$ can either be 0 or 1), then the target label is set to $1-y$. For multi-label datasets, since ImageNet contains 1,000 classes, the prediction and target labels differ for each dataset subset. Specifically, Cougar vs. Cheetah: predicted label is 286, corresponding to “cougar”; target label is 293, corresponding to "cheetah". We adopt this setting to perform a transition from the “cougar” class to the “cheetah” class. Sorrel vs. Zebra: The predicted label is 339, which corresponds to "sorrel"; the target label is 340, which corresponds to "zebra". We adopt this setting to perform a transition from the “sorrel” class to the “zebra” class. Egyptian Cat vs. Persian Cat: predicted label is 283, corresponding to "Persian cat"; target label is 285, corresponding to "Egyptian cat". We adopt this setting to perform a transition from the “Persian cat” class to the “Egyptian cat” class. For a detailed mapping between classes and their corresponding labels, please refer to this link.
Hyperparameters Settings We provide the detailed hyperparameter settings for each dataset in Table 6. $\alpha$ is the weight of the adversarial defense loss, $\lambda$ is the weight of the $l_{1}$ loss used during mask generation, and $\gamma$ is the threshold for generating the binary mask.
Table 6: Hyperparameter Settings for All Datasets.
Datasets $\alpha$ $\lambda$ $\gamma$ PGD Iterations Diffusion Steps
CelebA Smile 0.5 0.001 0.1 50 500
CelebA Age 0.5 0.001 0.1 50 500
CelebA-HQ Smile 0.05 0.001 0.1 50 500
CelebA-HQ Age 0.05 0.001 0.1 50 500
BDD-OIA 0.005 0.001 0.05 50 1000
Sorrel-Zebra 0.1 0.001 0.15 25 1000
Cougar-Cheetah 0.1 0.001 0.15 25 1000
Persian Cat-Egyptian Cat 0.05 0.001 0.15 25 1000
B. Additional Experimental Results
Results on CelebA-HQ Due to the similarity between the CelebA-HQ and CelebA datasets, we omit the CelebA-HQ results from the main text and provide them in this section as supplementary material. We first present the quantitative analysis. Tables 7 and 8 summarize the results on the CelebA-HQ dataset, which exhibit a similar overall trend to that observed on CelebA. Although DiME achieves slightly better FR, CD, and sFID scores, it does so at a significant expense of FID. Moreover, DiME is obviously less competitive in terms of perceptual similarity (LPIPS) and cosine similarity (S3). Notably, its COUT score is negative, indicating that the classifier lacks confidence in the counterfactual transitions produced by the DiME.
Table 7: Evaluation Results on CelebA-HQ: Smile.
Methods FR $\uparrow$ COUT $\uparrow$ LIPIPS $\downarrow$ S3 $\uparrow$ CD $\downarrow$ FID $\downarrow$ sFID $\downarrow$
DiVE - - - - - 107.0 -
STEEX - - - - - 21.9 -
DiME 90.00% -0.2797 0.1689 0.9245 2.52 15.8 26.79
ACE 77.50% 0.2369 0.0353 0.9868 2.76 5.9 38.20
Ours 89.35% 0.3918 0.0190 0.9892 2.74 4.4 27.85
Table 8: Evaluation Results on CelebA-HQ: Age
Methods FR $\uparrow$ COUT $\uparrow$ LIPIPS $\downarrow$ S3 $\uparrow$ CD $\downarrow$ FID $\downarrow$ sFID $\downarrow$
DiVE - - - - - 107.5 -
STEEX - - - - - 26.8 -
DiME 94.55% -0.3339 0.1689 0.9202 4.00 17.8 26.54
ACE 63.45% 0.0567 0.0448 0.9876 5.73 9.7 33.95
Ours 89.50% 0.3683 0.0278 0.9897 5.53 5.6 27.92
We then present the qualitative analysis. Since ACE achieves the strongest performance among all baselines, we focus our comparison on the counterfactual images generated by our method and those produced by ACE. As shown in Figure 7, similar to the results on the CelebA dataset, the counterfactual images generated by ACE contain many unintended modifications. These include changes in teeth and chin color, the appearance of exaggerated facial creases, and the emergence of extraneous patches on the cheeks. In the transition from “old” to “young,” the counterfactual images generated by our method exhibit comparable visual quality to those produced by ACE, with no obvious undesired changes in either case. As a result, we do not highlight any specific regions for this comparison. However, our method aligns more closely with the intended transformation—for example, by effectively removing age spots and smoothing wrinkles—demonstrating greater semantic consistency with the target concept.
More Results on Other Datasets The extended qualitative results on CelebA, BDD-OIA, and ImageNet are presented in Figures 8,  9, and  10, respectively.
Figure 7: Qualitative results on CelebA-HQ dataset. Red circles highlight unintended changes produced by ACE, while green circles indicate the corresponding improvements made by our method.
Figure 8: Extended qualitative results on CelebA dataset. Red circles highlight unintended changes produced by ACE, while green circles indicate the corresponding improvements made by our method.
Figure 9: Extended qualitative results on BDD-OIA dataset. Green circles highlight the counterfactual changes. The right column for each transition direction shows a zoomed-in version of the left column.
Figure 10: Extended qualitative results on ImageNet dataset. Green circles highlight the counterfactual changes.
C. VLM Prompt Design
In this section, we provide the prompt designs for each dataset to facilitate evaluation using Vision-Language Models (VLMs). While the evaluation dimensions are consistent with traditional metrics, the key distinction lies in our use of a VLM as a querying agent. Rather than relying solely on conventional metrics—which may be limited in capturing perceptual or semantic nuances—we engage the VLM through human-centric, dialog-style questions. This enables more holistic and perceptually grounded evaluations, aligning better with human judgment and interpretability.
CelebA Smile You are given two sets of facial images: one set contains the original images, and the other contains the corresponding counterfactual images. The target attribute to be modified is smile. If the person in the original image is smiling, the target is "not smiling"; if the person is not smiling, the target is "smiling". Please evaluate the counterfactual images based on the following four criteria, each rated on a scale from 1 to 10:
  • Validity: Does the counterfactual image successfully transition to the target attribute (i.e., smiling or not smiling) in a visually convincing way?
  • Sparsity: Are irrelevant attributes and regions left unchanged, with only the necessary parts being modified?
  • Proximity: From a human perceptual perspective, how close is the counterfactual image to the original one?
  • Realism: Does the counterfactual image remain close to the data manifold and distribution, i.e., does it appear as a realistic image that could exist in the original dataset?
CelebA Age You are given two sets of facial images: one set contains the original images, and the other contains the corresponding counterfactual images. The target attribute to be modified is age. If the person in the original image is old, the target is "young"; if the person is young, the target is "old". Please evaluate the counterfactual images based on the following four criteria, each rated on a scale from 1 to 10:
  • Validity: Does the counterfactual image successfully transition to the target attribute (i.e., smiling or not smiling) in a visually convincing way?
  • Sparsity: Are irrelevant attributes and regions left unchanged, with only the necessary parts being modified?
  • Proximity: From a human perceptual perspective, how close is the counterfactual image to the original one?
  • Realism: Does the counterfactual image remain close to the data manifold and distribution, i.e., does it appear as a realistic image that could exist in the original dataset?
CelebA-HQ Smile You are given two sets of facial images: one set contains the original images, and the other contains the corresponding counterfactual images. The target attribute to be modified is smile. If the person in the original image is smiling, the target is "not smiling"; if the person is not smiling, the target is "smiling". Please evaluate the counterfactual images based on the following four criteria, each rated on a scale from 1 to 10:
  • Validity: Does the counterfactual image successfully transition to the target attribute (i.e., smiling or not smiling) in a visually convincing way?
  • Sparsity: Are irrelevant attributes and regions left unchanged, with only the necessary parts being modified?
  • Proximity: From a human perceptual perspective, how close is the counterfactual image to the original one?
  • Realism: Does the counterfactual image remain close to the data manifold and distribution, i.e., does it appear as a realistic image that could exist in the original dataset?
CelebA-HQ Age You are given two sets of facial images: one set contains the original images, and the other contains the corresponding counterfactual images. The target attribute to be modified is age. If the person in the original image is old, the target is "young"; if the person is young, the target is "old". Please evaluate the counterfactual images based on the following four criteria, each rated on a scale from 1 to 10:
  • Validity: Does the counterfactual image successfully transition to the target attribute (i.e., smiling or not smiling) in a visually convincing way?
  • Sparsity: Are irrelevant attributes and regions left unchanged, with only the necessary parts being modified?
  • Proximity: From a human perceptual perspective, how close is the counterfactual image to the original one?
  • Realism: Does the counterfactual image remain close to the data manifold and distribution, i.e., does it appear as a realistic image that could exist in the original dataset?
BDD-OIA You are given two sets of animal images: one set contains the original images, and the other contains the corresponding counterfactual images. The target attribute is traffic status. If the traffic participants and infrastructure in the image indicate that the current traffic condition is stop or slow down, then the target is moving forward. Conversely, if the image reflects a moving forward traffic state, the target is to change it to stop or slow down. Please evaluate the counterfactual images based on the following four criteria, each rated on a scale from 1 to 10:
  • Validity: Does the counterfactual image successfully transition to the target attribute (i.e., smiling or not smiling) in a visually convincing way?
  • Sparsity: Are irrelevant attributes and regions left unchanged, with only the necessary parts being modified?
  • Proximity: From a human perceptual perspective, how close is the counterfactual image to the original one?
  • Realism: Does the counterfactual image remain close to the data manifold and distribution, i.e., does it appear as a realistic image that could exist in the original dataset?
Sorrel - Zebra You are given two sets of animal images: one set contains the original images, and the other contains the corresponding counterfactual images. The target attribute to be modified is animal type. The query attribute is "sorrel", and the target attribute is "zebra". Please evaluate the counterfactual images based on the following four criteria, each rated on a scale from 1 to 10:
  • Validity: Does the counterfactual image successfully transition to the target attribute (i.e., smiling or not smiling) in a visually convincing way?
  • Sparsity: Are irrelevant attributes and regions left unchanged, with only the necessary parts being modified?
  • Proximity: From a human perceptual perspective, how close is the counterfactual image to the original one?
  • Realism: Does the counterfactual image remain close to the data manifold and distribution, i.e., does it appear as a realistic image that could exist in the original dataset?
Cougar - Cheetah You are given two sets of animal images: one set contains the original images, and the other contains the corresponding counterfactual images. The target attribute to be modified is animal type. The query attribute is "cougar", and the target attribute is "cheetah". Please evaluate the counterfactual images based on the following four criteria, each rated on a scale from 1 to 10:
  • Validity: Does the counterfactual image successfully transition to the target attribute (i.e., smiling or not smiling) in a visually convincing way?
  • Sparsity: Are irrelevant attributes and regions left unchanged, with only the necessary parts being modified?
  • Proximity: From a human perceptual perspective, how close is the counterfactual image to the original one?
  • Realism: Does the counterfactual image remain close to the data manifold and distribution, i.e., does it appear as a realistic image that could exist in the original dataset?
Persian Cat - Egyptian Cat You are given two sets of animal images: one set contains the original images, and the other contains the corresponding counterfactual images. The target attribute to be modified is animal type. The query attribute is "Persian cat", and the target attribute is "Egyptian cat". Please evaluate the counterfactual images based on the following four criteria, each rated on a scale from 1 to 10:
  • Validity: Does the counterfactual image successfully transition to the target attribute (i.e., smiling or not smiling) in a visually convincing way?
  • Sparsity: Are irrelevant attributes and regions left unchanged, with only the necessary parts being modified?
  • Proximity: From a human perceptual perspective, how close is the counterfactual image to the original one?
  • Realism: Does the counterfactual image remain close to the data manifold and distribution, i.e., does it appear as a realistic image that could exist in the original dataset?
D. Additional Parameters Study
Except for the results on BDD-OIA, the patterns observed in the other datasets are consistent with those presented in the main text. For the BDD-OIA dataset, our method is able to achieve higher LPIPS and COUT scores while maintaining the FR without degradation when $\alpha=0.005$. Although the FR does not decrease within the range of $\alpha$ values considered in Figure 14, it is foreseeable that further increasing $\alpha$ would gradually lead to a decline in FR. On the BDD-OIA dataset, our method is unlikely to achieve a higher FR than ACE. Therefore, further increasing the value of $\alpha$ in pursuit of an ideal trade-off is no longer feasible. The setting $\alpha=0.005$ already yields satisfactory results, which is why we choose it as the final hyperparameter value. We focus on this parameter study—also serving as an ablation when $\alpha=0$—to examine the effect of $\alpha$, while keeping all other hyperparameters consistent with those used in ACE. This ensures a fair comparison and highlights the critical role of $\alpha$ in our method, as it controls the degree of causal constraint on the change of spurious factors.
Figure 11: Line plots of FR and FID with respect to varying values of $\alpha$ on CelebA Age.
Figure 12: Line plots of FR and FID with respect to varying values of $\alpha$ on CelebA-HQ Smile.
Figure 13: Line plots of FR and FID with respect to varying values of $\alpha$ on CelebA-HQ Age.
Figure 14: Line plots of FR and FID with respect to varying values of $\alpha$ on BDD-OIA.
Figure 15: Line plots of FR and FID with respect to varying values of $\alpha$ on ImageNet (Sorrel - Zebra).
Figure 16: Line plots of FR and FID with respect to varying values of $\alpha$ on ImageNet (Cougar - Cheetah.
Figure 17: Line plots of FR and FID with respect to varying values of $\alpha$ on ImageNet (Persian Cat - Egyptian Cat.